1. Calculate the population size (N) at the given time (t) with either r or R below: Calculate r or R from the given R or r aIf I could please have some guidlines on how to do this, I would greatly appreciate it! I'm totally lost!
Answer #1

Answer is provided below:

No = Initial population size t = time (generation) Nt = Population after generation t R = Net reproductive rate. r = Per capiN = Nox Ry 4252 = 100 x R25 R 25 - 42-52 R = 252 42.52 = 11.162] 25 ® t=3, No = 1000, Nt - 64 64 = 1000 x ert rx3 64 - = 0.06- 125 O.U 5 t=3, No = 1000 , R =0.5 Nt = 1000x (0.5) - 125 ert = 125 = 0.125 2x3 = en (0.125) = -2.08 ra - 2.08 -0.693 ) .03